dANIEL06:

A line passes through (3, –2) and (6, 2). Write an equation for the line in point-slope form. Rewrite the equation in standard form using integers. y + 2 = (x + 3); –4x + 3y = –18 y + 2 = (x – 3); –4x + 3y = –18 y – 2 = (x – 3); –4x + 3y = 18 y – 3 = (x + 2); –4x + 3y = 17

dANIEL06:

help

dANIEL06:

bro help

Florisalreadytaken:

hey i can just leave the thread, and continue doing my job instead of answering your little stupid quiz test so be patient okay :)

dANIEL06:

sorry just playing chiiillllll :)

dANIEL06:

back to topic

Florisalreadytaken:

I wasn't intending to reply since I don't have much time but since you need an answer this badly I might actually make some time Well we know that a point slope equation of a line consists of \(y-y_{1}=m\left(x-x_{1}\right)\) right? now, we're iven two coordinates: \( (3 \ , \ –2)\) & \((6\ ,\ 2)\) first hwe have to find the slope\((m)\) by using a formula which basically somes from the formula above, and that is \( \boxed{m=\dfrac{ (y-y_1) }{(x-x_1)}} \) That would leave us with \( m=\dfrac{ (2--2) }{(6-3)} =\boxed{\dfrac{ 4 }{3}}\) so, putting that into our first form of the formula (use any of the two points; I picked the second one because it's basically the same thing, and also to let you work with the \((3, –2)\) coordinate ) \(y-2=\dfrac{4}{3}(x-6)\) \(3(y-2)=4(x-6)\) \( 3y-6=4x-24 \) \( 3y=4x-18 \) \( \boxed{18 =4x-3y}\) that doesn't fit any of the options they just want to mess with you, so multiply it by one \( -1(18 =4x-3y)\) \(\boxed{ -18=-4x+3y }\) and I believe that is the answer